OpenStudy (marigirl):

For what values of d will the roots of 1/4x^2 +dx-8 =0 be real? Answer reads: All values Not sure what I am doing wrong, i thought b^2-4ac >0 for real values...

OpenStudy (anonymous):

You're on the right track. For this problem, your determinant is \[d ^{2}-4\times \frac{ 1 }{ 4 }\times(-8)\] or \[d ^{2}-8\] So you're \[d ^{2}\] must be greater than or equal to 8 in order to be positive and create real solutions.

OpenStudy (marigirl):

Hi, Thank you for taking the time out to answer my question.I substituted in values and got \[d ^{2}^{+8}\ge0\]

OpenStudy (marigirl):

\[d ^{2}+8\ge0\]

OpenStudy (marigirl):

I guess when you look at it...all numbers will satithatsatisfy that inequation

Directrix (directrix):

I was looking at this problem because it is interesting. I found the discriminant and got d² + 8. A value of d that makes d² + 8 < 0 would yield a nonReal root. d² + 8 would never be negative or zero. It appears that any d will do.

OpenStudy (anonymous):

Oops, I put minus instead of plus. Yep, you got it.

OpenStudy (marigirl):

That is why the answer reads All values,because all values will satisfy it. if we go on to solve it and find d, it becomes square root of a negative number ??? better to solve it by thinking about it?

Directrix (directrix):

I don't know your level of study but I think I can correctly say that the coefficients of the polynomials with which you will work for the time being will be real values. So, solving for d and getting a nonReal result would not be a value of d which applies to this problem. So, yes. I agree with you. Thinking about a problem conceptually is always appropriate and especially when interpreting an answer.